How much difference is there between your apparent weight when standing on flat ground at the equator v.s. at the south pole v.s. at 60 degree latitude? Assume Earth's radius is 6,380 km. What period of rotation for Earth would make you feel weightless at the equator?
Answer #1

equater, a=0; de = g- Rw²652 So, gequater = 9, = g - Rw265 0=g-Rw? g south = 92 = g_ Rw²6560 = g-RW² x 14 91-923 g - Rw*- g +